SUZAKU
Top
Suzaku
Suzaku Flash
Research
Research News
SWG
Tech Description
Presentation
Proposal
Accepted Targets
Schedule
Plan
Log
Process
Archive
Analysis
CALDB
Help Desck
HXD-WAM
Public Data
Bibliography
SWG
Japanese
Top > Suzaku > Log > HXD

HXD onboard event log

Updated on Apr. 30, 2014
(Changes in analysis software is not shown here.)

2014


6/12   08:46:00   HXD-S/AE-on (HXD events started to be output).

5/26   22:20:35   HXD-S/AE-0ff to save the power

5/16   11:57:00   HXD-on (HXD events started to be output).

5/15   20:33:00   HXD-S/AE/DE Off due to UVC

5/15   05:12:00   HXD-S/AE-on (HXD events started to be output).

5/12   10:51:00   HXD-S/AE-0ff to save the power

4/28   02:37:00   HXD-S/AE-on (HXD events started to be output).

4/24   18:49:00   HXD-S/AE-0ff to save the power

4/22   19:31:00   HXD-S/AE-on (HXD events started to be output).

4/15   23:27:00   HXD-S/AE-0ff to save the power

2/26   02:11:00   HXD-S/AE-on (HXD events started to be output).

2/25   00:10:00   HXD-S/AE-0ff to save the power

2/18   23:23:06   Set the appropriate command for HXD-PIN noise cut.
                Normal HXD data started to be output.

2/15   23:55:00   HXD-S/AE-on (HXD events started to be output).
                HXD-PIN noise cut does not work well.

2/12   02:00:00   HXD-DE on

1/09   18:39:19   HXD-OFF by UVC

1/03   22:49:05   PIN-LD at DE changed.
                (W08P1: 30->36, W09P1: 33->37, W10P0: 35->38).
                Processed data do not be affected by this change.

2012


July 23 
 20:05  HV-OFF AE-OFF (due to RCS operation)
 23:22  AE-on, HV-on, event normal output

July 13 
 00:16  AE-ON HV-PMT ON
 03:35  HV-PIN on, event normal output

July 11 
 21:00  HV-OFF AE-OFF (due to RCS operation)

April 27 (U_GEM)
 09:20:00  HV-P1 was down to 0V due to that W10-P0 became noisy since
 4/22 0:00.
 11:03     HV-P1 was again raised up to 400V.

March 26-27 (Crab)
 3/26 12:45:00 - 3/27 12:26:00  PINLD 4544@AE, PINLD at DE was set to be old.
 3/27 4:10:10-4:20:10 PMTUD off

March 23 (V1082_SGR)
 05:20;00-05:30:00  PINLD 4544@AE, PINLD at DE was set to be old.

March 15 (Crab)
 02:50:00-03:00:00  PINLD 5545@AE, PINLD at DE was set to be old.

Feb. 28-29 (Crab)
 2/28 11:08:00 - 2/29 13:23:00  PINLD 8777@AE
(at this time, the default PINLD at AE is 6656.)  

Jan 31  (just before GX_304-1 (406060010))
10:02:55   HXD was completely recovered, and normal HXD data started to be output.


Jan 30  
10:00:00   HXD-DE on. HXD data started to be output.
13:26:10   HXD-DE reset


Jan 24  (end of WKK_4438 (706011010))
01:27:46   HXD (S,AE,DE) became off due to UVC. 
           From this time to Jan 31, HXD data were not output.

2011



May 25 (from SMC_X-1 (706030050))
 16:11-     PIN lower threshold operation to reduce data saturation (both FIFO-full and 
             BUFFER FLUSH).
             Hereafter WPU-1 and 3 PIN triggering analog threshold was shifted from 0x05 to 0x06,
             to reduce both the data transfer saturation between HXD-AE and DE (FIFO-dull), 
             and the on-board telemetry saturation (BUFFER FLUSH).
             This operation affect all the 32 PIN diodes belonging to WPU-1 and 3 
             (PIN 16^31 and 48-63),
             and also affects the detector response at its lowest energy bandpass.
             Then, hereafter the new epoch 11 started.

March 21 (CRAB (105029010))
  March 21 19:21 - March 23 03:29
             PIN lower threshold of analog tragger of all PINs were raised for calibration.
             WPU0,1,2,3 ... LD=0x07,06,06,06 


2010


Dec 16 (from CygX-1(905006010))
  22:42-     The LD cut levels at DE (CPU) for most of PINs were
             raised. No effects for analysis.

April 5 (E0102-72)
  06:00-12:00 Every PMT-UD levels were temporarily increased to check the 
              suppression efficiency of triggers in analog electronics of HXD.
              GSO data during this period have cut-off at higher energies around 500 keV. 

April 3 (from the start of W28-CENTER (505005010))
              3rd PIN lower threshold operation to reduce data saturation.
              WPU-0 PIN triggering analog threshold was shifted from 0x05 to 0x06,
              to reduce both the data transfer saturation between HXD-AE and DE (FIFO-dull), 
              and the on-board telemetry saturation (BUFFER FLUSH).
              This operation affect all the 16 PIN diodes belonging to WPU-0 (PIN 0-15),
              and also affects the detector response at its lowest energy bandpass.

              3rd new detector response valid from April 3 will be provided by the HXD team.

Feb.2 (from the start of PERSEUS_1_4_WIN (104021010))
              Another PIN lower threshold operation to reduce data saturation (both FIFO-full and BUFFER FLUSH).
              WPU-2 PIN triggering analog threshold was shifted from 0x04 to 0x05,
              to reduce both the data transfer saturation between HXD-AE and DE (FIFO-dull), 
              and the on-board telemetry saturation (BUFFER FLUSH).
              This operation affect all the 16 PIN diodes belonging to WPU-2 (PIN 32-47),
              and also affects the detector response at its lowest energy bandpass.

              Another new detector response valid from Feb.2 (epoch 8) will be provided by the HXD team.

Jan.16 (from the start of 1H0419-577)
              PIN lower threshold operation to reduce data saturation (both FIFO-full and BUFFER FLUSH).
              Hereafter WPU-3 PIN triggering analog threshold was shifted from 0x04 to 0x05,
              to reduce both the data transfer saturation between HXD-AE and DE (FIFO-dull), 
              and the on-board telemetry saturation (BUFFER FLUSH).
              This operation affect all the 16 PIN diodes belonging to WPU-3 (PIN 48-63),
              and also affects the detector response at its lowest energy bandpass.

              In addition, W21-P1 in-orbit software lower cut threshold chage from 30(0x1e) to 33(0x21),
              to reduce on-board telemetry saturation (BUFFER FLUSH). Since this value is lower 
              than on-ground processing thershold, there is no effect on the cleaned data.
              
              New detector response valid from Jan. 16 to Feb. 2 (epoch 7) will be provided by the HXD team.

2009

Sep.13 (HESS_J1809-193B -- HESS_J1809-193E)
  00:20-18:45 Count rate of HXD W01p3 (PIN7) suddenly flares up. 
              This event is considered as those similar to flare-like events 
              events which previously happened in W10p0 and W01p3, the same one. 
              Event starts right after the passage of SAA. 
              Both lower and upper discriminators show increased rates.
              Event calmed down to normal rate spontaneously. 
              No additional operation took place.               

July.1 (FORNAX_A_W-LOBE_S)
  11:11-11:24 HXD-DE memory dump operation and reset and run (HK output).
  12:52-13:05 HXD-DE observation mode, HXD-AE parameter table upload.
  14:34-14:47 HXD-AE/DE parameter table upload. 
              Bias voltages of PIN diodes were increased to 100 V.
  16:15-16:29 High voltage of PMTs are increased to ~900 V (nominal).
              Bias voltages of PIN diodes were increased to 400/500 V (nominal0.
              Event data were output to the telemetry.
  17:57-18:09 Tuning of AE/DE parameters.
              The HXD is recovered to the nominal operation mode.

Jun.30 (A1795_FAR_SOUTHWEST/FORNAX_A_W-LOBE_S)
  11:13-11:25 HXD-DE reset and memory dump operation.
  12:54-13:07 HXD-DE memory dump operation.
  14:36-14:49 CPU run (HK telemetry output) and reset.

Jun.27 (A1795_NEAR_WEST)
  17:27- A sudden halt of HXD-DE occurred during a SAA passage.
         No telemetry data avarilable until the recovery operation on July.1.
         The bias voltage of all PIN diodes and PMTs are automatically 
         reduced to zero.

2007


01/27 (IRAS13224-3809)
  12:51-13:03, 14:33-14:45  
  The PMT gains for WAM units were temporarily changed during the earth occultation, to recover ~30% gain drop since the launch.

02/10 (E0102.2-7219)
  02/10 22:10- 02/11 19:20
  The PMT gains for WAM units were temporarily changed during the observation,
  to check the effect on the PIN background.

02/11 (E0102.2-7219)
  14:00-16:30, 16:38-18:04
  Events with the PIN-UD flag were output to the telemetry, for the purpose
  of the HXD-PIN background study.

03/18 (E0102.2-7219)
  03/18 21:00 - 03/19 07:10
  SUD events from W31 unit were output to the telemetry, for the purpose
  of the HXD-PIN background study.

03/20 (Crab)
  03/20 10:25 - 03/21 00:20
  The PMT gains for WAM units were temporarily changed during the observation,
  to check the effect on the PIN response.

04/10 (E0102.2-7219) 
  The PMT gains for WAM units were permanentlly changed.
  The GRB detection threshold was also changed by mistake, and a number of
  fake GRB triggers were detected until the trigger condition was fixed on 
  the next day.

06/08 (FAIRALL_9)
  The PMT UD level was changed from C9 to CA during the PSD-off period.

06/10 (47Tuc)
  The LD cut level for W21p1 inside the HXD-DE was increased to 30 from 25.

07/28 (NGC1386)
  Flare-like events from W10p0 were observed again after ~21:00.
  One of four high-voltage units (HV-P1) was reduced to 0V during
  07/30 16:45-18:30, and set at 300V during 07/30 23:35 - 07/31 16:50.

08/11 (IC4329A)
  The PMT UD level was permanently changed from 0xC9 to 0xCA,
  to reduce the data transfer rate between HXD-AE and HXD-DE.

 10:56:04 
 The PMT UD level was permanently changed from 0xC9 to 0xCA,
 to reduce the data transfer rate between HXD-AE and HXD-DE.
 PINLD of W10p0 was up to 32ch at DE.

08/29 (2S0921-630)
  The PSD function in HXD-AE and HXD-DE remained to be OFF after 
  06:10 by mistake. It was recovered to be ON at 07:09 in HXD-AE,
  while it remained to be OFF until 02:33 on 2007-08-30 in HXD-DE.
  The HXD data during this period suffered from huge dead times.

2006

1/4  (NGC720)
     11:37  PMT-HV recovered to 900V

3/14 (GX349+2 #1) 
      23:08-23:28 GSO PSD cut level tested (Fast=150ch -> Slow=150ch)
  	Shape of the GSO spectra around 40-100 keV changes during these priods.

3/15  00:45-01:10 GSO PSD cut level tested (PMTLD=0x04)
      02:25-02:50 GSO PSD cut level tested (Slow=140ch LD=0x04)
      04:00-04:20 GSO PSD cut level tested (Slow=140ch LD=0x05)
 	Shape of the GSO spectra around 40-100 keV changes during these priods.

3/19 (GX349+2 #2)
      23:06-23:24 GSO PSD cut level tested (Slow=130ch LD=0x04)
	Shape of the GSO spectra around 40-100 keV changes during these priods.

3/20  00:25  WAM time resolution changed from 1.0sec to 0.5sec
	Hereafter, the WAM-GRB data has two times higher time resolution,
	while the WAM-TRN data stays at the same.
      00:42-01:05 GSO PSD cut level tested (Slow=120ch LD=0x04)
	Shape of the GSO spectra around 40-100 keV changes during these priods.

3/22 (DRACO_HVC-B)
      00:07 HCE control temperature of the HXD sensor base plate 
	  changed from -14C to -20C. Basically no impact to science data.

3/23 (RXJ1856.5-3754)
      22:39 telemetry distribution table changed (Bit-M 33kbps -> 40kbps)
      22:51 GSO LD level changed from 0x05 to 0x04
      22:51-23:05 GSO PSD cut level tested (Slow=100ch)
      23:05-23:19 GSO PSD cut level tested (Slow=110ch)
      23:19-      GSO PSD cut level changed (Slow=120ch)
	Shape of the GSO spectra around 40-100 keV changes.

    Note: Later, increase of dead time in a high rate condition (i.e. after a few SAA passages
	and/or at low COR or high source count) is identified. This was caused by
	too high transer rate from HXD-AE to DE. Thus the GSO-LD level was changed
	back to 0x05 later in 5/13. Observations till 5/13 will be affected by a few short periods of 
	dead time increase around/after the SAA.

3/27 (GC#2)
     16:03- GSO PSD cut level changed (slow=100ch)
	Shape of the GSO spectra around 40-100 keV changes.

4/17 (PupA Interior, PupA_NE )
     13:56 Bit-M/L = 15kbps, GSO 4 units mode at Bit-M
	 Special operation to cope with the high XIS counting rate of the source.
4/18 07:25 Bit-M=40kbps, GSO 16 units mode at Bit-M
	 Back to normal operation after the special one.

4/19 (1E1841-045)
     10:33 GSO PSD cut level for W30 unit changed (Slow=120ch)
	Shape of the GSO spectra of unit W30 around 40-100 keV changes.

4/29 (1RXS_J213344.1+510)
     06:57 Bit-L data distribution table changed 10kbps --> 15kbps
	to avoid telemetry saturation at lower COR time region in Bit-L condition.

5/2  (4U1907+09)
     Command seqence was updated to acquire onboard monitoring histogram data of 
         PIN noise spectrum during PSD off dayly monitor operation (witihin earth occultation)

5/13 (SDSSJ084234.9+362)
      16:33 GSO LD level changed from 0x04 to 0x05 due to 
      the data transfer saturation between AE and DE.
      Observations from 3/23 to this moment are affected by the increase
      of dead time. See note on 3/23.

5/17 (LockmanHole)
      18:32-18:45 HXD-DE memory dump operation. Observation stops during this operation.
          PMT HV was set at 0V to allow the pipeline software to descard the data obtained wihtin this operation.

5/24 (1ES1959+650)
     ~10:00 Counting rate for one of 64 PIN (W01p3) gradually 
            increased from nominal rate (<50Hz) up to 300 Hz.
            The spectrum of this PIN showed a very hard index.
                Special care must be taken to analyze the data.

5/25 (CygnusLoop P15)
         Recovery operation starts.
     ~13:25 Bias from one of four high-voltage units (HV-P0) was set 
            down to 400 V (nominal = 500V), and the rate of W01p3 
            recovered to the nominal level. 
                Sixteen out of 64 PIN has different bias value after this operation.

5/26 (1ES1101-232)
         Recovery operation continues.
     ~11:40 Bias voltage of HV-P0 was increased up to 430 V.
     ~13:20 Bias voltage of HV-P0 was increased up to 490 V.
     The counting rate of W01p3 stayed at the nominal level.
     ~18:20 The rate of W01p3 again showed a gradual increase.
            Later it reached 930 Hz at the maximum.
                Note that the other PINs suffer no problem
                from the begining of this issue (at 5/24)


5/27 (BL Lac)
         Recovery operation continues.
     ~11:40 Bias voltage of HV-P0 was decreased down to 400V.
     ~13:20 Bias voltage of HV-P0 was decreased down to 0V.
            PIN Events from four well units (W00,01,02,03) were
            triggered off.
                Sunday starts and the HXD is operated with a quarter of its PINs off
                for ~1.5 days. Note that the other PINs suffer no problem
                from the begining of this issue (at 5/24)


5/28 (CH Cyg) 
5/29 (Cygnus Loop-1) Bias voltage of HV-P0 was kept at 0V.

5/29 (Cygnus Loop-2)
     ~11:40 Bias voltage of HV-P0 was increased up to 200V. 
     ~13:25 Bias voltage of HV-P0 was increased up to 400V.
            PIN events from W0x units are again enabled to output.
     ~15:05 Bias voltage of HV-P0 was increased up to 460V.
     ~16:40 Bias voltage of HV-P0 was decreased down to 400V.

     This bias has been kept hereafter. 1/4 of the PINs are 
     operated with a bias of 400 V, and 3/4 with 500 V.
     Although the change in response and background are not large (~<10%),
     the detailed study is underway (6/25 2006--).
     Note that the other PINs suffer no problem
     from the begining of this issue (at 5/24)

6/16 (1ES1426+428)
     17:30 The gains of three WAM units (T13,14,33) were changed 
     so as to be the same as those of other units.
     This operation in principle affects the hit-pattern mask, 
     but at present(6/25 2006), no difference is detected in the data of the Well-units.


9/20 (IGRJ16195-4945)
  ~0:00 Counting rate for one of 64 PIN (W10p0) starts showing a few small flare-like events.
  The count rate increase is modest (~1Hz this day and <10 Hz at 2nd Oct.).
  It was invisible in the HK light curve (raw trigger rate), and was detected only after screening. 
  Thus the phenomena was not recognized untill 2nd October.
  The number and intensity of the flare-like events slowly increases hereafter.
  Spectra of the flares are very soft, and phenomenologicaly it looks as an increase in the noise level.
  	   
  Data of the observations after this target should be handled with care. 

10/03 (ABELL_2199_OFFSET_2)
  ~ 23:35 Operation for W10P0 flare-like event starts.
  Three out of the four PIN-HV was set to 0V. The other one (WPU0-PIN) is still set at 400 V.
  Therefore, only 16 out of 64 PINs are active. In summary,
  WPU0/1/2/3 PIN-HV are changed from 400/500/500/500 V to 400/0/0/0 V until 2006-10-04 23:30.

10/26 (ABELL_3376_WEST_RELIC)
  11:12-11:22 HXD-DE memory dump operation. Observation stops during this operation.
  PMT HV was set at 0V to allow the pipeline software to descard the data obtained wihtin this operation.

12/30 (NGC 4649) 
  High-voltage of HXD was set to zero at 2006-12-30 05:04. Returned to
  normal value at 2007-01-04 03:16. No event data during this period.

2005

7/23 HXD-DE ON
7/25 HXD-AE ON
7/26-8/5  HV operation / parameter adjustment
8/15-8/17 HV operation cont.
8/19 (CenA) First-light

8/20 (A2052 #1-4) trial of "GSO 4 units mode", becomes a Bit-Low standard operation
     PASS-1 PIN software LD channel changed (+3-4 ch)

8/22 (Crab) boresight calibration
     PASS-1 HXD_TIME_LATCH (the first time of time counter clear)
            GRB data output enabled (WAM)
            W30 hit-pattern analog LD level changed (0x35->0x25)

8/23 PASS-1 PIN onboard software LD channel changed (+1 ch)

8/24 (PSR B1509) RBM was activated (PMT-HV reset to 0V)
	Data is not useful around ...
     	2005/08/24 11:30:37 - 2005/08/24 11:46:33
      	2005/08/24 13:01:33 - 2005/08/24 19:17:17

8/27 PASS-2 RBM threshold level changed (T0,1 5->10kHz)
8/27 07:30:00 - 8/28 03:50:00 (Bulge)
		Hit pattern width changed (0x0A->0x0F) for test
		PIN background level is different during this period

8/28 03:50:00 - Hit pattern width changed (0x0F->0x0A)
		PIN BGD level changes

8/30 (N103B) RBM was activated
	Data is not useful around ...
        2005/08/30 06:27:05 - 2005/08/30 08:30:07

*9/2 (Vega) Hit-pattern width changed to 0x0F (PIN background is reduced)
		Fixed to this value hereafter. 
		PIN BGD level changes.

9/3 (NEP) RBM was activated
	Data is not useful around ...
        2005/09/03 06:27:05 - 2005/09/03 06:44:57

9/4 (SN1006)
    19:37 - 9/7 17:51 event data stopped since OP stopped. 
	This is due to ground station operation cancel caused by typhoon 

9/9 (SN1006_NE_RIM) RBM was activated (due to high solar activity)
	Data is not useful around ...
        2005/09/09 09:55:33 - 2005/09/09 12:20:01
        2005/09/09 19:40:01 - 2005/09/09 20:42:35
        2005/09/10 02:58:49 - 2005/09/10 03:17:33

9/11 (SN1006_SW_RIM) RBM was activated
	Data is not useful around ...
        2005/09/11 07:20:53 - 2005/09/11 14:00:53

9/12 (SN1006_SW_BG) RBM was activated
	Data is not useful around ...
        2005/09/12 08:46:25 - 2005/09/12 10:35:29
        2005/09/12 12:57:13 - 2005/09/12 14:00:01

9/13 (FORNAX) RBM was activated
	Data is not useful around ...
        2005/09/13 19:23:11 - 2005/09/13 20:35:55
        2005/09/13 23:18:25 - 2005/09/13 23:52:25

9/15 (CrabHXD, CrabXIS)
     16:40 - 22:50 Hit pattern width changed (0x0F->0x0A) for test
		Note that this change is temporal only for calibration.
		0x0F is the default.

9/17 (NGC2110)
     14:20 RBM level changed (T0,1=100kHz, T2,3=10kHz) to decrease 
	 the frequency of RBM activation by chance. 
	 Basically no impact to science data. 

10/22 (Mkn3)
      01:53-01:57 HXD-DE memory dump (temporal observation stop)
      01:57       W30 hit-pattern LD level changed (0x25->0x35)
      03:34-03:38 HXD-DE error clear (temporal observation stop)
	 
10/30 (AE_AQUARII)
      HXD_TIME_LATCH command at the time of each Manuevor (MNV) has been changed 
	  from non-clear to clear mode (the second time of time counter clear). 
	  Hereafter, at every MNV the HXD time counters are cleared.
	  On-source observation data with calibrated time assignment has no difference.
      (Only a few seconds of data in the MNV mode may have mis-assigend time.)
	  
11/10 (NGC4051)
	The definition of the SAA region is revised.
	Hereafter, the timing of high-voltage reduction is slightly shifted,
	and the duration is shortened (optimized).
	This new version of SAA definition is applied for all observations
	in the screening criteria of the data process. Note that the criteria is also designed 
	to properly handle both the data obtained before and after this revision.

11/16 (A1413)
      08:41 HCE control temperature on the HXD sensor base-plate 
	changed from -20C to -14C. Basically no impact to science data.

12/15 (SafeHold)
      01:51-02:04 GSO PSD cut level changed (Fast=150ch -> Slow=150ch)
      03:27-03:46 GSO PSD cut level changed (Slow=150ch -> 120ch)
	this change was temporal.

12/30 (NGC720)
      12:19- PMT-HV set to 0V (new year operation)
	  	Data is not useful

ISAS/JAXA Department of High Energy Astrophysics

Last Modified: Tuesday, 10-Jun-2014 17:17:06 JST